# Investigating and Defending Shortcut Learning in Personalized Diffusion Models

### v3

@@ Abstract
Personalized diffusion models have gained popularity for adapting pre-trained text-to-image models to generate images of specific topics with minimal training data. However, these models are vulnerable to minor adversarial perturbations, leading to degraded performance on corrupted datasets. Such vulnerabilities are further exploited to craft protective perturbations on sensitive images like portraits that prevent unauthorized generation. In response, diffusion-based purification methods have been proposed to remove these perturbations and retain generation performance. However, existing works lack detailed analysis of the fundamental shortcut learning vulnerability of personalized diffusion models and also turn to over-purifying the images, which causes information loss. In this paper, we take a closer look at the fine-tuning process of personalized diffusion models through the lens of shortcut learning. And we propose a hypothesis explaining the manipulation mechanisms of existing perturbation methods, demonstrating that perturbed images significantly deviate from their original prompts in the CLIP-based latent space. This misalignment during fine-tuning causes models to associate noisy patterns with identifiers, resulting in performance degradation. Based on these insights, we introduce a systematic approach to maintain training performance through purification. Our method first purifies the images to realign them with their original semantic meanings in latent space. Then, we introduce contrastive learning with negative tokens to decouple the learning of clean identities from noisy patterns, which shows a strong potential capacity against adaptive perturbation. Our study uncovers shortcut learning vulnerabilities in personalized diffusion models and provides a firm evaluation framework for future protective perturbation research. Code is available at https://github.com/liuyixin-louis/DiffShortcut.
@@ 1 Introduction
The rapid advancements in text-to-image diffusion models, such as DALL-E 2 [1], Stable Diffusion [2], and MidJourney [3], have revolutionized the field of image generation. These models can generate highly realistic and diverse images based on textual descriptions, enabling a wide range of applications in creative industries, entertainment, and beyond. However, the ability to fine-tune these models for personalized generation using a small set of reference images has raised concerns about their potential misuse, such as generating misleading or harmful content targeting individuals [4, 5] or threatening the livelihood of artists by mimicking unique artistic styles without compensation [6].
@@
To address these issues, several data-poisoning-based methods have been proposed to protect user images from unauthorized personalized synthesis. These methods aim to proactively make images resistant to AI-based manipulation by crafting adversarial perturbations [5, 7], applying subtle style-transfer cloaks [6], or crafting misleading perturbation that causes model’s overfitting [4, 8]. The model trained on perturbed data will generate images poor in quality, and thus, the unauthorized fine-training fails. Despite the protection effectiveness, different from the adversarial perturbation crafted for fixed off-the-shelf diffusion models, where the protection against unauthorized editing [7] can be well explained by the adversarial vulnerability of neural networks [9], and also the sharpness of the latent space of VAE [10, 11]. The underlying mechanism for why protective perturbation disturbs the fine-tuning of the diffusion model has not been explored yet.
@@
Furthermore, purification studies are also purposed to further break those protections. As demonstrated in [4], most of the protection methods lack resilience against minor data transformations like Gaussian smoothing or JPEG compression. However, this simple rule-based purification causes severe data quality degradation. As demonstrated in [12, 13], diffusion-reconstructed-based purification shows a strong capacity to denoise the images and yield high-quality output by leveraging the distribution modeling ability of diffusion models. Based on the observation that clean images have better consistency upon LDMs-based reconstruction, IMPRESS [12] proposes optimization on the images to ensure reconstruction consistency with visual LPIPS similarity constraints. Despite effectiveness, such purification requires a tremendous of time due to the iterative nature of the propose optimization. On the other line, DiffPure [14] proposes to leverage off-the-shelf pixel-space diffusion models to conduct an SDEdit process that converts the perturbed images into a certain noisy state with a diffusion forward process and then denoises in the reverse process. GrIDPure [15] further divides the images into smaller grids and employs small-step DiffPure that yields better visual consistency. However, these SDEdit-based purification approaches still yield unfaithful content that causes great change in identity due to the generative nature of the diffusion model leveraged. How to design an effective, efficient, and faithful purification approach is still an open question.
@@
Given the above two major limitations of existing works, in this paper, we first take a closer look at the reason behind this protection, where we found the latent of the perturbed images by these methods are largely shifted from its original concept. Going deeper, we discover that the diffusion models fine-tuned on these perturbed images turn to link the rarely-appeared concept token with the injected noisy perturbation instead of learning the clean identity shown in images. Based on these findings, we showcase the shortcut-learning vulnerability of personalized diffusion upon fine-tuning on poison.
Motivated by this empirical understanding, we present a systematic purification and training approach to empower robust personalized diffusion models upon perturbed data. Our approach conducts comprehensive purification from three perspectives, including input image-text pair purification, contrastive learning with the negative token, and quality-enhanced sampling. Compared to existing purification methods that are only limited to image denoising, the advantages, and benefits of our framework are three folds: i) efficiency and faithfulness: we conduct image purification by using off-the-shelf super-resolution and image restoration models that convert low-quality, noisy images into high-quality, purified ones; ii) robustness and once-for-all: we demonstrate that contrastive learning with negative token shows strong ability in defending against adaptive attack crafted with knowledge on the purification models; iii) systematic: not limiting to image denoising, we conduct a systematic investigation on other potential approaches that a model trainer can leverage in training and sampling, to better facilitate the evaluate the comprehensive evaluation on the effectiveness and resistance of future protection methods. Our main contributions are summarized as follows:
  • We empirically explain the underlying mechanism of the success of recent protective perturbations for safeguarding against unauthorized personalize diffusion model fine-tuning, where we find that the perturbed images no longer match the prompt pair in latent space, and thus it misleads the models on learning the wrong concept.
  • Based on this understanding, we propose a systematic defensive framework that empowers robust personalized diffusion models from input image-text pair, model training, and the sampling process. Specifically, on the input denoising, we leverage off-the-shelf image restoration models to realign the image with its semantic meaning in the prompt; on model training, we propose contrastive learning with the negative token that leverages the clean class-prior dataset to decouple the learning of false relationship further; on the sampling, we design simple quality-enhanced approach with negative prompting. As a result, our method is more effective, efficient, and faithful.
  • We empirically demonstrate the effectiveness and faithfulness of our framework on the facial dataset VGGFace2 and painting dataset WikiArt. Moreover, we demonstrate the robustness of our defending framework under adaptive attack when other variants fail.
@@ 2 Related Works
Latent Diffusion Models and Customized Adaption.
Generative Models (GMs) aim to synthesize samples from a data distribution given a set of training examples. Diffusion Probabilistic Models [16] are now the dominant GMs with various applications such as text-to-image synthesis [17], image editing [18, 19, 20], and image inpainting [21]. To improve the training efficiency, Latent Diffusion Models (LDMs) [2] are proposed to conduct a more efficient diffusion process in a low-dimensional latent space with a pair of pre-trained image encoder and decoder (e.g., VAE [10]). Despite the general capability, model users often wish to synthesize specific concepts from their own personal lives. To meet that, customized text-to-image generation [22, 23, 24, 25] is proposed to learn the concepts using a small set of reference images. Among them, DreamBooth [23] is one of the mainstream approaches that involve full fine-tuning of parameters, yielding more superior generation performance compared to those only fine-tune partial parameter [26] or a pseudo-word vector [27]. Besides, it uses a rarely-used token to link the concept and incorporate prior-preserving loss with class-oriented images, enabling a high-quality and diverse generation of subjects.
Data Poisoning as Protection against Unauthorized Data Usage with LDMs.
@@
Recent text-to-image LDMs have raised concerns about their potential misuse in generating misleading or harmful content targeting individuals [4, 5] and threatening the livelihood of professional artists by style mimicking [6]. To address these issues, several data-poisoning-based methods have been proposed to protect user images from unauthorized personalized synthesis by injecting adversarial perturbations through minimizing adversarial target loss in image encoder or UNet denoiser [6, 5], or denoising-loss maximization [7, 4, 8] or in opposite direction, denoising-loss minimization [28], or cross-attention loss maximization [29]. Despite its effectiveness, the underlying mechanism of protection against diffusion model fine-tuning has not been well explored yet. To the best of our knowledge, [13] is the only work that attempts to investigate this. However, [13] only shows some relevance of the vulnerability to the text encoder. In this work, we take a closer look at this mechanism and provide a more comprehensive explanation from a latent mismatch and shortcut learning perspective.
Data Purification that Further Breaks Protection.
@@
Despite promising protection performance, studies [4, 8, 13] suggest that these perturbations are brittle and can be easily removed with minor transformations. However, adaptive protection with EoT [30] indicates that protection can further bypass these simple rule-based transformations. Another direction is to leverage diffusion models as purifiers to perturb images back to their clean distributions. In the classification scenario, DiffPure [14] is a mainstream approach for adversarial purification by applying SDEdit [31] on the poison with an off-the-shelf diffusion model. For purification against protective perturbation, GrIDPure [32] further adapts iterative DiffPure with small steps on multi-grid spitted image to perverse the original resolution and structure. However, due to their generative nature, these SDEdit-based purifications have limitations in yielding unfaithful content, where the purified images fail to preserve the original identity. Observing the perceptible inconsistency between the perturbed images and the diffusion-reconstructed ones, IMPRESS [12] conducts the purification via minimizing the consistency loss with constraints on the maximum LPIPS-based [33] similarity change on pixel space. While it manages to preserve similarity, IMPRESS suffers from the inefficiency issue due to its iterative process and is ineffective under stronger protections like [8].
@@ 3 Preliminary
Latent Diffusion Models (LDMs). Generative diffusion model [16] aims to model an unknown data distribution by learning an iterative denoiser to map random variables from pre-defined Gaussian distribution to those real data samples. Latent diffusion model [2] further conducts such a process in a lower-dimensional latent space to improve training and inference efficiency. Furthermore, by training on prompt-image pair, LDMs can take additional embedding as conditions to guide the denoising process, enabling it to generate or edit images with guidance from a prompt (e.g., a text). In LDMs, an image encoder $\mathcal{E}$ is first used to map the image $\mathbf{x}_{0}$ into a latent representation, i.e., $\mathbf{z}_{0}=\mathcal{E}(\mathbf{x}_{0})$. And then LDMs craft noisy latent $z_{t}$ in forward phrase with $q_{t}(z_{t}\mid z_{0})=\mathcal{N}(z_{t};\sqrt{\bar{\alpha}_{t}}\,z_{t-1},\,(1-\bar{\alpha}_{t})\mathbf{I})$, where $\beta_{t}$ growing from 00 to $1$ are pre-defined values, $\alpha_{t}=1-\beta_{t}$, and $\bar{\alpha}_{t}=\Pi_{s=1}^{t}\alpha_{s}$. Suppose we have a textual prompt $c$, and a text encoder $\tau_{\theta}$ that yield the embedding $\bm{c}=\tau_{\theta}(c)$, The goal of the LDM is to train a conditional noise estimator network $\bm{\epsilon}_{\theta}$, e.g., a UNet [34], by predicting the Gaussian noise added in previous timestamp, to model the conditional distribution $p(\mathbf{z}_{0}|\bm{c})$ by gradually recovering $\mathbf{z}_{0}$ from $\mathbf{z}_{T}$ with additional textual information $c$. Suppose $\bm{\epsilon}_{\theta}(\mathbf{z}_{t},t,\bm{c})$ is the Gaussian noise estimated in the $t$-th step and $\bm{\epsilon}$ is the grouth-truth Gaussian noise sampled for $\mathbf{z}_{t-1}$. On each image-text pair $(\mathbf{x}_{0},\bm{c})$, the denoising network $\bm{\epsilon}_{\theta}$ is trained based on the following loss:
$\mathcal{L}_{\text{denoise}}(\mathbf{x}_{0},\bm{c};\theta):=\mathbb{E}_{\mathbf{z}_{0}\sim E(\mathbf{x}_{0}),\bm{\epsilon},t}\Big{[}\|\bm{\epsilon}-\bm{\epsilon}_{\theta}(\mathbf{z}_{0},t,\bm{c})\|_{2}^{2}\Big{]}.$ (1)
During the inference phrase, we first sample $\mathbf{z}_{T}\sim\mathcal{N}(0,\textbf{I})$ and then conduct multi-steps denoising to obtain a latent $\tilde{\mathbf{z}}_{0}$ with potential prompt conditioning $\bm{c}$, and then we can obtain generated image with an image decoder $\mathcal{D}$, i.e., $\tilde{\mathbf{x}}_{0}=\mathcal{D}(\tilde{\mathbf{z}}_{0})$.
Personalized LDMs via Dreambooth Fine-tuning. DreamBooth [23] is a fine-tuning-based method aimed at updating LDMs for generating images of specific concepts. The main idea of Dreambooth is to introduce a new identifier to link the subject concepts and use a class-specific prior-preserving loss to mitigate overfitting and language-drifting issues. Specifically, it fine-tunes a pre-trained LDM on two additional text-image sets, including an instance dataset $\mathcal{D}_{\bm{x_{0}}}=\left\{\left(\bm{x_{0}}^{i},\bm{c}^{\text{ID}}\right)\right\}_{i}$ and a class dataset $\mathcal{D}_{\bar{\bm{x}_{0}}}=\left\{\left(\bar{\bm{x}}_{0}^{i},\bm{c}\right)\right\}_{i}$, where the subject image is $\bm{x}_{0}$, and the class image is $\bar{\bm{x}}_{0}$. For the prompt design, the class-specific prompt $C$ is set as “a photo of a [class noun]" and the instance prompt $C^{\text{ID}}$ is set as “a photo of sks [class noun]", where sks specifies the subject and “[class noun]" is the category of object (e.g., “person"). Formally, DreamBooth first generates prior class data using the original checkpoint and then optimizes a weighted sum of instance denoising loss and prior-preservation loss,
@@
where $\lambda$ controls for the relative importance of this term. With approximately $\sim 1000$ training steps and $\sim 4$ subject images, it can generate vivid personalized subject images with Stable Diffusion [35].
Protective Perturbation against Personalized LDMs. Recent studies [4] suggest that minor adversarial perturbation to clean images can disturb the learning of customized diffusion and also prevent image editing with an off-the-shelf diffusion model by greatly degrading the quality of the generated image. Existing protective perturbation can be classified into two categories: perturbation crafted with fixed diffusion models and perturbation crafted with noise-model alternative updating. In this paper, we focus on the second category since they are more effective in the fine-tuning setting. The general framework of these protective perturbation methods is to craft noise that maximizes an adversarial loss $\mathcal{L}_{adv}$ that is typically designed as the denoising loss $\mathcal{L}_{denoise}$ and also alternatively update the noise generator surrogates $\bm{\theta}^{\prime}$ can be a single model [4] or an ensemble of models [8]). Furthermore, the surrogates might only update part of their components [29]. Formally, at the $j$-th alternative step, the noise surrogate $\bm{\theta}^{\prime}_{j}$ and perturbation $\bm{\delta}^{(j)}$ are updated via solving,
$\small\,\,{\bm{\theta}^{\prime}}_{j}\leftarrow\underset{{{\bm{\theta}}^{\prime}}_{j-1}}{\mathrm{arg}\min}\sum_{x}{\mathcal{L}_{db}}\left(\mathbf{x}+\bm{\delta}^{(j-1)},\bm{c}^{ID},\bar{\bm{x}},\bm{c};{{\bm{\theta}}^{\prime}}_{j-1}\right);\bm{\delta}^{(j)}\leftarrow\underset{\left\lVert\bm{\delta}^{(j-1)}\right\rVert_{\infty}\leq r}{\mathrm{arg}\max}\mathcal{L}_{adv\,\,}\left(\mathbf{x}+\bm{\delta}^{(j-1)},\bm{c};{{\bm{\theta}}^{\prime}}_{j}\right).$ (3)
To solve this, standard Gradient Descent is performed on the model parameter while the images are updated via Project Gradient Descent (PGD) [36] to satisfy the $\ell_{\infty}$-ball perturbation budget constrain,

### v4

# Rethinking and Defending Protective Perturbation in Personalized Diffusion Models

@@ Abstract
Personalized diffusion models (PDMs) have become prominent for adapting pretrained text-to-image models to generate images of specific subjects using minimal training data. However, PDMs are susceptible to minor adversarial perturbations, leading to significant degradation when fine-tuned on corrupted datasets. These vulnerabilities are exploited to create protective perturbations that prevent unauthorized image generation. Existing purification methods attempt to mitigate this issue but often over-purify images, resulting in information loss. In this work, we conduct an in-depth analysis of the fine-tuning process of PDMs through the lens of shortcut learning. We hypothesize and empirically demonstrate that adversarial perturbations induce a latent-space misalignment between images and their text prompts in the CLIP embedding space. This misalignment causes the model to erroneously associate noisy patterns with unique identifiers during fine-tuning, resulting in poor generalization. Based on these insights, we propose a systematic defense framework that includes data purification and contrastive decoupling learning. We first employ off-the-shelf image restoration techniques to realign images with their original semantic meanings in latent space. Then, we introduce contrastive decoupling learning with noise tokens to decouple the learning of personalized concepts from spurious noise patterns. Our study not only uncovers fundamental shortcut learning vulnerabilities in PDMs but also provides a comprehensive evaluation framework for developing stronger protection. Our extensive evaluation demonstrates its superiority over existing purification methods and stronger robustness against adaptive perturbation. Please find our code at https://github.com/liuyixin-louis/DiffShortcut.
@@ 1 Introduction
The rapid advancements in text-to-image diffusion models, such as DALL-E 2 (Ramesh et al., 2022), Stable Diffusion (Rombach et al., 2022), and MidJourney (mid, ), have revolutionized the field of image generation. These models can generate highly realistic and diverse images based on textual descriptions, enabling a wide range of applications in creative industries, entertainment, and beyond. However, the capability to fine-tune these models for personalized generation using a small set of reference images has raised concerns about their potential misuse, such as generating misleading or harmful content targeting individuals (Van Le et al., 2023; Salman et al., 2023) or threatening the livelihood of artists by mimicking unique artistic styles without compensation (Shan et al., 2023).
@@
To address these issues, several data-poisoning-based methods have been proposed to protect user images from unauthorized personalized synthesis (Šarčević et al., 2024; Deng et al., 2024; Wang et al., 2024). These methods aim to proactively make images resistant to AI-based manipulation by crafting adversarial perturbations (Salman et al., 2023; Liang et al., 2023), applying subtle style-transfer cloaks (Shan et al., 2023), or crafting misleading perturbation that causes model’s overfitting (Liu et al., 2024). The model trained on perturbed data will generate images that are poor in quality, and thus, the unauthorized fine-tuning fails. Despite the protection effectiveness, different from the adversarial perturbation crafted for fixed off-the-shelf diffusion models, where the protection against unauthorized editing (Liang et al., 2023) can be well explained by the adversarial vulnerability of neural networks (Ilyas et al., 2019), and the sharpness of the latent space of VAE (Kingma & Welling, 2013; Guo et al., 2023; Xue et al., 2023). The underlying mechanism for how protective perturbation disturbs the fine-tuning of the personalized diffusion model has not been explored yet.
@@
Moreover, purification studies are also purposed to further break those protections. As demonstrated in Van Le et al. (2023), most of the protection methods lack resilience against minor data transformations like Gaussian smoothing or JPEG compression. However, this simple rule-based purification causes severe data quality degradation. As demonstrated in Cao et al. (2024); Zhao et al. (2024), diffusion-reconstructed-based purification shows a strong capacity to denoise the images and yield high-quality output by leveraging the distribution modeling ability of diffusion models. Based on the observation that clean images have better consistency upon reconstruction, IMPRESS (Cao et al., 2024) proposes optimization on the images to ensure reconstruction consistency with visual LPIPS similarity constraints (Zhang et al., 2018). Despite effectiveness, such purification requires a tremendous amount of time due to the iterative nature of the proposed optimization. On the other line, DiffPure (Nie et al., 2022) proposes to leverage off-the-shelf pixel-space diffusion models to conduct an SDEdit process (Meng et al., 2021) that converts the perturbed images into a certain noisy state with a diffusion forward process and then denoises in the reverse process. GrIDPure (Zheng et al., 2023) further divides the images into smaller grids and employs small-step DiffPure that yields better visual consistency. However, these SDEdit-based purification approaches still yield unfaithful content that causes great change in identity due to the generative nature of the diffusion model leveraged. How to design an effective, efficient, and faithful purification approach is still an open question.
@@
To tackle these limitations, we first take a closer look at the fine-tuning process of PDMs through the lens of shortcut learning(Geirhos et al., 2020). We hypothesize that adversarial perturbations introduce a latent-space misalignment between images and their textual prompts in the latent space. This misalignment causes the model to erroneously associate the injected noise patterns with the unique identifier tokens during fine-tuning, resulting in degraded performance. To validate our hypothesis, we conduct comprehensive experiments demonstrating that perturbed images significantly deviate from their semantic concepts in the latent space. Fine-tuning on such mismatched image-prompt pairs leads the model to learn spurious correlations toward noise patterns.
Based on these insights, we propose a systematic defense framework to empower robust PDMs against adversarial perturbations. Our approach conducts comprehensive purification from three perspectives, including input image purification, contrastive decoupling learning with the negative token, and quality-enhanced sampling. Compared to existing purification methods that are only limited to image purification, the advantages of our framework are three folds: i) efficiency and faithfulness: we conduct image purification by using off-the-shelf super-resolution and image restoration models that convert low-quality, noisy images into high-quality, purified ones; ii) robustness and once-for-all: we demonstrate that contrastive decoupling learning itself alone works and contributes in defending against adaptive attacks crafted against the purification pipeline; iii) system-level defense: not only limited to the input image, we propose systematic defending strategies covering three stages including data purification, model training, and sampling strategy, offering a comprehensive evaluation on the effectiveness of future protection methods. Our main contributions are summarized as follows:
  • We uncover and empirically validate the shortcut learning vulnerabilities in PDMs caused by latent-space image-prompt misalignment due to adversarial perturbations.
  • We propose a systematic defense framework based on causal analysis that effectively mitigates these vulnerabilities through data purification and contrastive decoupled learning and sampling.
  • We demonstrate the effectiveness, efficiency, and faithfulness of our approach through extensive experiments across 7 protections, showing significant improvements over existing defenses. Our study provides a more robust evaluation framework for future research on protective perturbations.
@@ 2 Related Works
Data Poisoning as Protection against Unauthorized Training with LDMs. Latent Diffusion Models (LDMs) (Rombach et al., 2022) have become dominant in various generative tasks, including text-to-image synthesis. To meet the demand for personalized generation, methods like DreamBooth (Ruiz et al., 2023) have been proposed, which fine-tune LDMs using a small set of reference images to learn specific concepts. However, these advancements have raised concerns about potential misuse, such as generating misleading content targeting individuals (Van Le et al., 2023; Salman et al., 2023) and threatening the livelihood of professional artists through style mimicking (Shan et al., 2023). To address these issues, several data-poisoning-based methods have been proposed to protect user images from unauthorized personalized synthesis by injecting adversarial perturbations through minimizing adversarial target loss in image encoder or UNet denoiser (Salman et al., 2023), or denoising-loss maximization (Liang et al., 2023; Van Le et al., 2023; Liu et al., 2024) or in opposite direction, denoising-loss minimization (Xue et al., 2023), or cross-attention loss maximization (Xu et al., 2024). Despite its effectiveness, the underlying mechanism of protection against diffusion model fine-tuning has not yet been explored well. To the best of our knowledge, Zhao et al. (2024) is the only work that attempts to investigate the underlying mechanism. However, it is only limited to the vulnerability of the text encoder. In this work, we provide a more comprehensive explanation from the view of latent mismatch and shortcut learning.
@@
Data Purification that Further Breaks Protection. Despite promising protection performance, studies (Van Le et al., 2023; An et al., 2024; Liu et al., 2024) suggest that these perturbations without advanced transformation loss (Athalye et al., 2018) are brittle and can be easily removed under simple rule-based transformations. Among all types of transformation, state-of-the-art adversarial purification leverages diffusion models as purifiers to perturb images back to their clean distributions. In the classification scenario, DiffPure (Nie et al., 2022) is a mainstream approach for adversarial purification by applying SDEdit (Meng et al., 2021) on the poison with an off-the-shelf diffusion model. For purification against protective perturbation, GrIDPure (Lee & Chang, 2022) further adapts iterative DiffPure with small steps on multi-grid spitted image to preserve the original resolution and structure. However, due to their generative nature, these SDEdit-based purifications have limitations in yielding unfaithful content, where the purified images fail to preserve the original identity. Observing the perceptible inconsistency between the perturbed images and the diffusion-reconstructed ones, IMPRESS (Cao et al., 2024) conducts the purification via minimizing the consistency loss with constraints on the maximum LPIPS-based (Zhang et al., 2018) similarity change on pixel space. While it manages to preserve similarity, IMPRESS suffers from the inefficiency issue due to its iterative process and is ineffective under stronger protections like Liu et al. (2024); Mi et al. (2024).
Shortcut Learning and Causal Analysis. Shortcut learning occurs when models exploit spurious correlations in training data, leading to poor generalization (Geirhos et al., 2020). The causal analysis provides a framework for addressing this by modeling cause-effect relationships (Pearl, 2009; Schölkopf et al., 2021). It helps identify true causal factors, distinguishing them from spurious correlations. In computer vision, models may incorrectly focus on background textures instead of object features (Brendel & Bethge, 2019). Techniques like Invariant Risk Minimization (Arjovsky et al., 2019) and Counterfactual Data Augmentation (Teney et al., 2021) leverage causal principles to improve robustness. In PDMs, adversarial perturbations can introduce spurious correlations between noise patterns and identifiers during fine-tuning. Our work explores how to restore correct causal relationships when learning PDMs on perturbed data, which is under-explored in existing works.
@@ 3 Preliminary
Personalized Latent Diffusion Models via DreamBooth Fine-tuning. Latent Diffusion Models (LDMs) (Rombach et al., 2022) are generative models that perform diffusion processes in a lower-dimensional latent space, enhancing training and inference efficiency compared to pixel-space diffusion models (Ho et al., 2020). By conditioning on additional embeddings such as text prompts, LDMs can generate or edit images guided by these prompts. Specifically, an image encoder $\mathcal{E}$ maps an image $\mathbf{x}_{0}$ to a latent representation $\mathbf{z}_{0}=\mathcal{E}(\mathbf{x}_{0})$. A text encoder $\tau_{\theta}$ produces a text embedding $\bm{c}=\tau_{\theta}(c)$ for a given prompt $c$. The model trains a conditional noise estimator $\bm{\epsilon}_{\theta}$, typically a UNet (Ronneberger et al., 2015), to predict the Gaussian noise added at each timestep $t$, using the loss:
$\mathcal{L}_{\text{denoise}}(\mathbf{x}_{0},\bm{c};\theta)=\mathbb{E}_{\mathbf{z}_{0}\sim\mathcal{E}(\mathbf{x}_{0}),\bm{\epsilon},t}\Big{[}\|\bm{\epsilon}-\bm{\epsilon}_{\theta}(\mathbf{z}_{0},t,\bm{c})\|_{2}^{2}\Big{]}.$ (1)
During inference, the model starts from random noise $\mathbf{z}_{T}\sim\mathcal{N}(0,\textbf{I})$ and iteratively denoises it to obtain a latent $\tilde{\mathbf{z}}_{0}$, which is then decoded to generate the image $\tilde{\mathbf{x}}_{0}=\mathcal{D}(\tilde{\mathbf{z}}_{0})$. DreamBooth (Ruiz et al., 2023) fine-tunes a pre-trained LDM to generate images of specific concepts by introducing a unique identifier that links subject concepts and employing a class-specific prior-preserving loss to mitigate overfitting and language drift. The fine-tuning utilizes an instance dataset $\mathcal{D}_{\bm{x}_{0}}=\left\{\left(\bm{x}_{0}^{i},\bm{c}^{\text{ID}}\right)\right\}_{i}$ and a class dataset $\mathcal{D}_{\bar{\bm{x}}_{0}}=\left\{\left(\bar{\bm{x}}_{0}^{i},\bm{c}\right)\right\}_{i}$, where $\bm{x}_{0}$ are subject images and $\bar{\bm{x}}_{0}$ are class images. The class-specific prompt $C$ is set as “a photo of a [class noun]”, and the instance prompt $C^{\text{ID}}$ is “a photo of sks [class noun]”, where sks specifies the subject and “[class noun]” denotes the object category (e.g., “person”). The fine-tuning process optimizes a weighted sum of the instance denoising loss and the prior-preservation loss:
@@
where $\lambda$ balances the two terms. With approximately 1k training steps and around four subject images, DreamBooth can generate vivid, personalized subject images (von Platen et al., 2022). Protective Perturbation against Personalized LDMs. Recent studies suggest that minor adversarial perturbation to clean images can disturb the learning of customized diffusion and also prevent image editing with an off-the-shelf diffusion model by greatly degrading the quality of the generated image. Existing protective perturbation can be classified into two categories: perturbation crafted with fixed diffusion models and perturbation crafted with noise-model alternative updating. In this paper, we focus on the second category since they are more effective in the fine-tuning setting. The general framework of these protective perturbation methods is to craft noise that maximizes an adversarial loss $\mathcal{L}_{adv}$ that is typically designed as the denoising loss $\mathcal{L}_{denoise}$ and also alternatively update the noise generator surrogates $\bm{\theta}^{\prime}$ can be a single model (Van Le et al., 2023) or an ensemble of models (Liu et al., 2024)) or the attention modules (Xu et al., 2024). Formally, at the $j$-th alternative step, the noise surrogate $\bm{\theta}^{\prime}_{j}$ and perturbation $\bm{\delta}^{(j)}$ are updated via solving,
$\,\,{\bm{\theta}^{\prime}}_{j}\leftarrow\underset{{{\bm{\theta}}^{\prime}}_{j-1}}{\mathrm{arg}\min}\sum_{x}{\mathcal{L}_{db}}\left(\mathbf{x}+\bm{\delta}^{(j-1)},\bm{c}^{ID},\bar{\bm{x}},\bm{c};{{\bm{\theta}}^{\prime}}_{j-1}\right);\bm{\delta}^{(j)}\leftarrow\underset{\left\lVert\bm{\delta}^{(j-1)}\right\rVert_{\infty}\leq r}{\mathrm{arg}\max}\mathcal{L}_{adv\,\,}\left(\mathbf{x}+\bm{\delta}^{(j-1)},\bm{c};{{\bm{\theta}}^{\prime}}_{j}\right).\vspace{-2pt}$ (3)
To solve this, standard Gradient Descent is performed on the model parameter while the images are updated via Project Gradient Descent (PGD) (Madry et al., 2018) to satisfy the $\ell_{\infty}$-ball perturbation budget constrain,
@@
Causal Analysis and Structural Causal Model. Causal analysis models cause-and-effect relationships between variables (Pearl, 2009), helping identify spurious correlations and mitigate shortcut learning Geirhos et al. (2020). A Structural Causal Model (SCM) uses structural equations and a directed acyclic graph to represent causal relationships. It comprises endogenous variables $\mathbf{V}$, exogenous variables $\mathbf{U}$, and structural equations ${f_{i}}$, where each $V_{i}\in\mathbf{V}$ is defined as $V_{i}=f_{i}(\text{Pa}(V_{i}),U_{i})$. By intervening on spurious correlations, causal analysis helps models focus on true causal relationships rather than superficial patterns. For more details, see Pearl (2009); Geirhos et al. (2020).